OpenStudy (bekkah323):

4-1/x^2 / 2-1/x *equation below

OpenStudy (bekkah323):

\[\frac{ 4-\frac{ 1 }{ x^2 } }{ 2-\frac{ 1 }{ x } }\]

OpenStudy (anonymous):